# On Achievable Spectral Efficiency Using Adaptive Transmission Over Terrestrial Coherent FSO Links

### v9

@@ III-B Immediate Insights
The proposed ASE formula is interesting in obtaining quantitative as well as qualitative insights as follows.
  • At high SNRs, contribution from the Meijer-G based term vanishes while $\mu\to 1/\mathrm{SNR}$. Keeping this in mind, and now comparing the ASE of channels with and without pointing error, the ASE loss due to pointing error can be easily quantified in terms of $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ parameters.
  • Likewise, the ASE loss due to both turbulence and pointing error at high SNR when compared with no channel impairments (pure AWGN) can also be quantified and analyzed.
  • In a fixed terrestrial FSO link, variation in turbulence can lead to change in the pointing error behavior. In such a scenario, the derived closed-form ASE expression also allows an accurate analysis of the impact of these interlinked channel parameters.
A more in-depth and broader exposition of adaptive transmission benefits and the impact of channel impairments will be presented in Section IV, allowing for a wide range of variations in channel turbulence, pointing error, and SNR.
@@ III-C Low-SNR and High-SNR Asymptotic Expressions
[…]
@@
The threshold value $\mu$ associated with the optimal beam power adaptation as described in (14) for the transmitted laser beam satisfies
$\mathrm{(i)}\,\,\,\text{At high SNR:}\,\,\mu\,\approx\,\dfrac{1}{\mathrm{SNR}},$ (72)
$\mathrm{(ii)}\,\,\text{At low SNR:}\,\,\,\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\dfrac{1}{\mathrm{SNR}}\right).$ (73)
@@
Part (i): Substituting (9) into the average optical power constraint (15), we have
$\dfrac{\mu\mathrm{SNR}}{A}=\int_{\mu}^{\infty}\biggl(1-\dfrac{\mu}{\lambda}\biggr)\,G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-3.50006pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)d\lambda,$
where $A\,:=\,{ab\xi^{2}/(A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)})$. Applying the identity in (20) to the power constraint equation gives
$\dfrac{\mu\mathrm{SNR}}{A}\,=\,\dfrac{1}{A}\,$-\, $\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)$ (76)
$+\,\,\mu\,$ $G_{2,4}^{3,1}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1,0\end{array}\hskip-3.50006pt\right).$ (79)
From the power constraint in (15), it follows that $\mu$ exhibits a strictly monotonic inverse relation with $\mathrm{SNR}$, decreasing as $\mathrm{SNR}$ increases. Notably, at high SNRs, the contributions of the Meijer-G function terms in (76) become negligible due to their vanishing behavior for small arguments. Hence limμ→0    μSNR   =   1. This completes the proof of 1).
[…]
$G_{1,3}^{3,0}\left(z\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)\,\approx\,\dfrac{\sqrt{\pi}\,z^{(-7+2a+2b)/4}}{e^{2\sqrt{z}}}.$ (82)
Only the dominant term of the series expansion is shown in (82), and substituting it into the LHS of the constraint (15) yields
$\mathrm{SNR}\,$ $\approx\,\mathcal{K}\left(\dfrac{B_{1}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})}{ab\mu/A_{\mathrm{\scriptscriptstyle 0}}}-B_{2}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})\right)$ (83)
@@
$\Gamma(s,t)\,\approx\,t^{s}e^{-t}\biggl(\,\dfrac{1}{t}\,+\,\dfrac{s-1}{t^{2}}\,+\,o\left(\dfrac{1}{t}\right)^{3}\,\biggr).$ (86)
@@
$\mathrm{SNR}\,\approx\,\mathcal{K}\left(\dfrac{ab\mu}{A_{\mathrm{\scriptscriptstyle 0}}}\right)^{\left(\frac{a+b}{2}-\frac{11}{4}\right)}e^{-2\sqrt{\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu}}.$ (87)
@@
$\ln(\mathrm{SNR})\,\approx\,-2\sqrt{{ab\mu}/{A_{\mathrm{\scriptscriptstyle 0}}}}.$ (88)
@@
$\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\frac{1}{\mathrm{SNR}}\right).\vskip-10.00002pt$ (89)
This completes the proof of 2). ∎
@@
For the coherent HD optical channel subjected to GG turbulence with pointing errors, the asymptotic ASE at low and high SNR are given by
@@
where the superscripts $\mathrm{low}$’ and $\mathrm{high}$’ stand for the low-SNR (asymptotically 0) and high-SNR (asymptotically $\infty$) regimes respectively.
@@
Part (i): At high SNR, the threshold behaves as $\mu\approx\mathrm{SNR}^{-1}$ (from Lemma 1), while the Meijer’s G function becomes negligible for small input arguments, asymptotically approaching zero as
@@
$\left[\ln(z)-\psi(a)-\psi(b)+\dfrac{1}{\xi^{2}}\right].$ (94)
@@
Finally, substituting $\mu$ from  (89) in (96) completes the proof of the low SNR ASE expansion (90) in Theorem 2. ∎
A few interesting observations from the derived theorems are as follows.
  • By applying $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the asymptotic ASEs of the GG turbulence channel ‘without pointing errors’ at low and high SNR are computed as
@@
  • By comparing (91) and (98), a measure of the ASE degradation due to pointing errors at high SNRs can be found as
@@
The derived asymptotic results in (90), (91), (97) and (98) are useful for a comprehensive assessment of the impact of variations of turbulence and pointing error statistics over a wide range.
@@ IV Numerical Results and Discussion
These settings cover the practical range of ATP capabilities and provide a unified framework for analyzing turbulence-limited, jitter-limited, and intermediate regimes.
@@
$\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ (100)
which underscores its dependence on the index of refraction structure parameter $C_{\mathrm{n}}^{2}$, as discussed below.
@@
For near-ground horizontal propagation, the parameter $C_{\mathrm{n}}^{2}$ remains relatively constant and primarily varies with altitude-being stronger near the ground and weaker at higher elevations [11, Sec. 12.2][48, Ch. 2].
@@
  • Recall from (14) and (15) that $\mathrm{SNR}:=P_{\mathrm{avg}}/N_{\mathrm{hd}}$ is the average transmit SNR in all the derived ASE formulae, and that the horizontal axis in Figures 1-7 is the transmit SNR in log scale, not the received SNR. This enables fair comparison and accurate analysis mentioned earlier. If required, the average received SNR can be computed by scaling the transmit $\mathrm{SNR}$ by the mean channel gain $\mathbb{E}[\lambda]$.
@@
1) Exact ASE behavior: Based on Theorem 1, the exact ASE for HD-detection based coherent optical channel under GG turbulence, with and without pointing errors, is computed and shown in Figures 1 and 2, respectively. The analytical results closely match Monte Carlo simulation outcomes across a wide SNR range and under various turbulence strengths and pointing jitter conditions.
Figure 1: Average spectral efficiency of the GG turbulence channel without pointing errors under the coherent HD detection scheme. The average received SNR is the same as the transmit SNR since the mean channel gain $\mathbb{E}[\lambda]\,=\,1$.
Figure 2: Average spectral efficiency of the GG turbulence channel with pointing jitter under the coherent HD detection scheme. The received SNR is less than the transmit SNR by the mean channel gain $\mathbb{E}[\lambda]$.
Discussion: Figures 1 and 2 provide the following key observations:
  • As shown in Fig. 1, spectral efficiency deteriorates with increasing turbulence at high SNR. This is expected, as stronger turbulence leads to a heavier concentration of low channel gains in the PDF (cf. Fig. 6).
  • As shown in Fig. 2, a significant drop in spectral efficiency at high SNRs is observed in the presence of pointing errors. This is due to the effective SNR reduction at the receiver by the factor $\mathbb{E}[\lambda]$: milder jitter levels result in a moderate reduction in effective SNR, while severe jitter leads to a significant degradation in system performance.
  • Interestingly, Fig. 2 reveals that, under strong pointing errors, spectral efficiency improves with turbulence in the high-SNR regime: the transition from ASE degradation to improvement initiates at transmit SNR of $30$ dB, and correspondingly to a received SNR of $(30-17.58)\approx 12.42$ dB (since $\mathbb{E}[\lambda]=-17.58$ dB). To explain this counterintuitive behavior, an asymptotic high-SNR analysis is carried out next.
2) ASE behavior at High SNRs: The high-SNR asymptotic ASE behavior in (91), as established in Theorem 2, invites comparison with the ASE of an ideal AWGN channel, which is asymptotically given by $\ln\mathrm{SNR}$. The resulting difference—being strictly negative—quantifies the spectral efficiency loss at high SNRs due to the combined impairments of GG turbulence and pointing jitter, and is given by
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,\triangleq\,\ln\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)\,+\,\psi(a)\,+\,\psi(b)\,-\,\left(\frac{1}{\xi^{2}}\right).$ (101)
@@
Figure 3: The average spectral efficiency loss (in magnitude) at high SNRs as a function of turbulence strength for the GG channel with pointing jitter under the coherent HD detection scheme.
As shown in Fig. 3, for the GG fading with strong pointing error channel setting, the ASE at high SNR increases with turbulence in the practical $0<\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\leq 4$ range. This atypical ASE behavior with turbulence arises because the reduction in the pointing jitter effect more than compensates for the loss in collected power at the receiver. Consequently, the ASE penalty (magnitude) decreases approximately from $10$ to $7.1$ bits/s/Hz, indicating a remarkable improvement of nearly $3.0$ bits/s/Hz in the high SNR regime.
For the GG fading with mild pointing error channel setting, no such trade-off is observed, as increased $\xi$ induces marginal improvement while $A_{\mathrm{\scriptscriptstyle 0}}$ continues to decay significantly, leading overall to a monotonic ASE loss with turbulence. The corresponding high-SNR penalty magnitude, illustrated by the blue curve in Fig. 3, ranges from $3.2$ to $6.1$ bits/s/Hz.
In the GG fading without pointing error, on the other hand, the penalty at high SNRs is described by (see (98))
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,=\,\psi(a)\,+\,\psi(b)\,-\,\ln\left(ab\right)$ (102)
which corresponds to a relatively small reduction, as can be inferred from Fig. 2 (further qualification in Fig. 4).
Figure 4: Average spectral efficiency of the GG turbulence channel without pointing error under the coherent HD scheme: $A_{\mathrm{\scriptscriptstyle 0}}=1$ for “ideal” system condition whereas $A_{\mathrm{\scriptscriptstyle 0}}$ from Table I are considered as representation for realistic system condition.
@@
  • A common inconsistency in the literature is the omission of the parameter $A_{\mathrm{\scriptscriptstyle 0}}$ when modeling optical fading without pointing errors. While $A_{\mathrm{\scriptscriptstyle 0}}=1$ is often assumed for simplicity, practical FSO systems may exhibit $A_{\mathrm{\scriptscriptstyle 0}}<1$ due to beam divergence and finite aperture effects. Both cases are realistic in practice, but neglecting $A_{\mathrm{\scriptscriptstyle 0}}<1$, when not physically justified, can lead to an overestimated performance, as illustrated in Fig. 4.
@@
  • For the fixed-length optical channel with a constant FSO system configuration, as considered here, an important observation arises regarding the impact of increasing pointing jitter. Recall that $\sigma_{\mathrm{e}}$ represents the standard deviation of the pointing jitter. At high SNR, the corresponding ASE degradation (in nats/s/Hz) attributable solely to jitter can be obtained directly from (91) and is expressed as
@@
Figure 5: Spectral efficiency at low SNRs for the GG turbulence channel without pointing jitter under the HD detection scheme (see Table I for turbulence settings).
@@
$\text{Relative pre-factor}\,\approx\,\frac{a_{j}b_{j}}{a_{i}b_{i}}.$ (104)
Figure 6: The PDF of the GG fading gain for the settings described in Table I. Notice the larger probabilities for the higher gains with increasing turbulence strength.
Figure 7: Average spectral efficiency results in the low SNR regime for the GG turbulence channel with mild pointing jitter under the coherent HD detection scheme.

### v10

@@ III-B Immediate Insights
The proposed ASE limit is interesting in obtaining quantitative as well as qualitative insights as follows.
  • At high SNRs, contribution from the Meijer-G based term vanishes while $\mu\to 1/\rho$. Keeping this in mind, and now comparing the ASE of channels with and without pointing error, the ASE loss due to pointing jitters can be easily quantified in terms of $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ parameters.
  • Likewise, the ASE loss at high SNRs due to both turbulence and pointing error when compared with no channel impairments (pure AWGN) can also be quantified.
  • In a given terrestrial FSO link configuration, variations in turbulence statistics can influence the associated pointing error characteristics. The derived closed-form ASE expression enables accurate assessment of the coupled impact of these interdependent channel parameters.
A more in-depth and broader exposition of the impact of optical channel impairments and adaptive transmission benefits will be presented in Section IV, allowing for a wide range of variations in turbulence, pointing error, and SNR.
@@ III-C Low-SNR and High-SNR Asymptotic Expressions
The asymptotic expansions of the exact ASE expression should further reveal how key fading channel parameters from the GG turbulence and the pointing error distributions influence the performance of adaptive FSO schemes. To this end, the $\mu$$\rho$ relationship is first derived under average power constraint in both low-SNR and high-SNR regimes.
@@
The threshold $\mu$ corresponding to the optimal beam power adaptation in (15) satisfies
$\mathrm{(i)}\,\,\,\text{At high SNR:}\,\,\mu\,\approx\,\dfrac{1}{\rho},$ (73)
$\mathrm{(ii)}\,\,\text{At low SNR:}\,\,\,\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\dfrac{1}{\rho}\right).$ (74)
@@
Part $(i)$: Substituting (9) into the average optical transmit power constraint (16), we have
$\dfrac{\rho\mu}{\mathcal{B}}=\int_{\mu}^{\infty}\biggl(1-\dfrac{\mu}{\lambda}\biggr)\,G_{1,3}^{3,0}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\lambda\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-4.49997pt\right)d\lambda\,$,
where $\mathcal{B}:={ab\xi^{2}/(A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)})$. Applying the identity in (21) to the above power constraint equation gives
$\dfrac{\rho\mu}{\mathcal{B}}\,=\,\dfrac{1}{\mathcal{B}}\,$-\, $\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)$ (77)
$+\,\,\mu\,$ $G_{2,4}^{3,1}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1,0\end{array}\hskip-4.49997pt\right).$ (80)
From the power constraint in (16), it follows that $\mu$ exhibits a strictly monotonic inverse relation with $\rho$, decreasing as $\rho$ increases. Notably, at high SNRs, the contributions of the Meijer-G function terms in (77) become negligible due to their vanishing behavior for small arguments. Hence limμ→  0     ρμ  =   1. This completes the proof of (i).
Part (ii): We next derive an explicit $\mu$$\rho$ relationship in the low-SNR regime. Since the cutoff $\mu$ increases as $\rho$ decreases, power allocation is restricted to large channel gains. To this end, the channel distribution $f_{\lambda}(\lambda)$ in (9) is approximated using a low-order series expansion of the Meijer $G$-function for large arguments, as detailed below:
$G_{1,3}^{3,0}\left(z\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)\,\approx\,\dfrac{\sqrt{\pi}\,z^{(-7+2a+2b)/4}}{e^{2\sqrt{z}}}\,\cdot$ (83)
Only the dominant term of the series expansion is shown in (83), and substituting it into the LHS of (16) yields
$\rho\,$ $\approx\,\mathcal{K}\left(\dfrac{B_{1}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})}{ab\mu/A_{\mathrm{\scriptscriptstyle 0}}}-B_{2}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})\right)\,\mathpunct{\raisebox{2.15277pt}{,}}$ (84)
@@
$\Gamma(s,t)\,\approx\,t^{s}e^{-t}[\,{t}^{-1}\,+\,{(s-1)}{t^{-2}}\,+\,O({t}^{-3})].$ (87)
@@
$\rho\,\approx\,\mathcal{K}\left({ab\mu}/{A_{\mathrm{\scriptscriptstyle 0}}}\right)^{\left({a+b}-\frac{11}{2}\right)\frac{1}{2}}e^{-2\sqrt{{ab\mu}/{A_{\mathrm{\scriptscriptstyle 0}}}}}.$ (88)
@@
$\ln\rho\,\approx\,-2\sqrt{{ab\mu}/{A_{\mathrm{\scriptscriptstyle 0}}}}.$ (89)
@@
$\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\frac{1}{\rho}\right)\,\cdot\vskip-10.00002pt$ (90)
This completes the proof of (ii). ∎
@@
For the adaptive coherent terrestrial FSO channel in (10), with intensity gain $I=|h|^{2}$ distributed according to (9) and perfect channel state information (CSI) at both the transmitter and receiver, the asymptotic ASEs at low and high SNR are given by
@@
where the superscripts $\mathrm{low}$ and $\mathrm{high}$ correspond to the limits $\rho\to 0$ and $\rho\to\infty$, respectively.
@@
Part (i): At high SNR, the threshold scales as $\mu\approx\rho^{-1}$ (cf. Lemma 1), while the Meijer-G term vanishes for small arguments, i.e.,
@@
$\,\biggl[\,\ln z-\psi(a)-\psi(b)+\dfrac{1}{\xi^{2}}\,\biggr]\cdot$ (95)
@@
Finally, substituting $\mu$ from  (90) in (97) completes the proof of the low-SNR ASE expansion (91) in Theorem 2. ∎
Furthermore, we can deduce a few interesting observations from the derived theorems as follows.
  • By applying $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the asymptotic ASEs of the GG turbulence channel ‘without’ pointing errors at low and high SNRs are computed as
@@
  • The high-SNR ASE degradation solely due to pointing errors, obtained by comparing (92) and (99), is given by
@@
The asymptotic results in (91), (92), (98), and (99) enable comprehensive assessment of turbulence and pointing error effects across diverse terrestrial coherent FSO conditions.
@@ IV Numerical Results and Discussion
These FSO link settings cover the practical range of ATP capabilities and provide a unified framework for analyzing turbulence-limited, jitter-limited, and intermediate regimes.
@@
$\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}},$ (101)
which underscores its dependence on the index-of-refraction structure parameter $C_{\mathrm{n}}^{2}$, as discussed below.
@@
For a near-ground horizontal propagation path, the index-of-refraction structure parameter $C_{\mathrm{n}}^{2}$ remains unchanged and varies with height above ground [12, Section 12.2]. A wide range of optical turbulence conditions is achievable at different altitudes through $C_{\mathrm{n}}^{2}$; being stronger near the ground and weaker at higher elevations [45, Ch. 2].
@@
  • The horizontal axis in Figures 1-8 represents $\rho$ (in dB scale), i.e., the normalized transmit power, rather than the received SNR, ensuring a fair and consistent comparison. Recall from (12) that $\rho:=P_{\mathrm{avg}}/\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}$ is the noise-normalized transmit power budget in all the derived ASE formulae. If required, the average received SNR can be computed by scaling $\rho$ with the mean channel gain $\mathbb{E}[\lambda]$.
@@
1) Exact ASE behavior: Based on Theorem 1, the exact ASE for HD-detection based coherent optical channel under GG turbulence, with and without pointing errors, is computed and shown in Figures 1 and 2, respectively. The analytical results closely match Monte Carlo simulation outcomes across a wide SNR range and under various atmospheric turbulence regimes and pointing jitter conditions.
Figure 1: Average spectral efficiency achievable over a GG turbulence channel without pointing errors under adaptive coherent (I/Q) FSO transmission and synchronous heterodyne detection at the receiver. The average received SNR is the same as $\rho$ since the mean channel gain $\mathbb{E}[\lambda]\,=\,1$.
Figure 2: Average spectral efficiency achievable over a GG turbulence channel with pointing jitter under adaptive coherent FSO transmission and synchronous heterodyne detection at the receiver. The received SNR is less than $\rho$ by a factor equal to the mean channel gain $\mathbb{E}[\lambda]$.
Discussion: The following key observations are drawn from Figures 1 and 2:
  • ASE deteriorates with increasing turbulence at high SNR, as shown in Fig. 1; the PDF concentrates more heavily at lower channel gains as turbulence increases (cf. Fig. 6).
  • As shown in Fig. 2, ASE drops significantly at high SNRs in the presence of pointing errors. This is due to the effective SNR reduction at the receiver by the factor $\mathbb{E}[\lambda]$: milder jitter levels result in a moderate reduction in effective SNR, while severe jitter leads to a significant degradation in system performance.
  • Interestingly, Fig. 2 reveals that, under strong pointing errors, ASE improves with turbulence in the high-SNR regime: the transition from ASE loss to gain initiates at norm. tx. power $\rho=30$ dB, and correspondingly to a received SNR of $(30-17.58)\approx 12.42$ dB (since $\mathbb{E}[\lambda]=-17.58$ dB). To explain this counterintuitive behavior, an asymptotic high-SNR analysis is carried out next.
2) ASE behavior at High SNRs: The high-SNR asymptotic ASE behavior in (92), as established in Theorem 2, invites comparison with the ASE of an ideal AWGN channel, which is asymptotically given by $\ln\rho$. The resulting difference—being strictly negative—quantifies the spectral efficiency degradation at high SNRs due to the combined GG turbulence and pointing jitter impairments, and is given by
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,\triangleq\,\ln\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)\,+\,\psi(a)\,+\,\psi(b)\,-\,\left(\frac{1}{\xi^{2}}\right)\cdot$ (102)
@@
Figure 3: ASE degradation at high SNRs versus turbulence strength in a GG turbulence channel, for mild and strong pointing jitter, under adaptive coherent FSO transmission and synchronous heterodyne detection at the receiver.
As shown in Fig. 3, for the GG fading with ‘strong’ pointing errors, the ASE at high SNR increases with turbulence over the practical range $0<\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\leq 4$. This atypical behavior arises because the reduction in the pointing jitter effect more than compensates for the loss in collected power at the receiver. Consequently, the ASE penalty decreases approximately from 10-10 to 7.2-7.2 bits/s/Hz, yielding a gain of nearly $2.8$ bits/s/Hz in the high SNR regime.
For the GG fading with ‘mild’ pointing errors setting, no such trade-off is observed, as increased $\xi$ induces marginal improvement while $A_{\mathrm{\scriptscriptstyle 0}}$ continues to decay significantly, leading overall to a monotonic ASE loss with turbulence. The corresponding high-SNR penalty, illustrated by the purple curve in Fig. 3, ranges from 3.2-3.2 to 6.1-6.1 bits/s/Hz.
In the GG fading without pointing errors, on the other hand, the penalty at high SNRs is described by (see (99))
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,=\,\psi(a)\,+\,\psi(b)\,-\,\ln\left(ab\right),$ (103)
which results in a small ASE degradation, as can be inferred from Fig. 2 (Fig. 4 provides a more refined characterization).
Figure 4: Average spectral efficiency over a GG turbulence channel without pointing errors under adaptive coherent FSO transmission and synchronous heterodyne detection: $A_{\mathrm{\scriptscriptstyle 0}}=1$ corresponds to channel condition with full-beam capture, whereas the values of $A_{\mathrm{\scriptscriptstyle 0}}$ listed in Table I represent channel conditions with partial-beam capture. The definition of $A_{\mathrm{\scriptscriptstyle 0}}$ is given in (7).
@@
  • A common inconsistency in the literature is the omission of the parameter $A_{\mathrm{\scriptscriptstyle 0}}$ when modeling optical fading without pointing errors. $A_{\mathrm{\scriptscriptstyle 0}}=1$ suggests full beam capture which is completely justifiable for short-haul optical links where beam divergence is small and receive aperture is wide enough to capture the whole beam. $A_{\mathrm{\scriptscriptstyle 0}}<1$ implies partial capture due to strong beam divergence and finite aperture effects. Both cases are realistic in practice, but neglecting $A_{\mathrm{\scriptscriptstyle 0}}<1$, when not physically justified, can lead to an overestimated performance, as illustrated in Fig. 4.
@@
  • For the fixed-length optical channel with a constant FSO system configuration, as considered here, an important observation arises regarding the impact of increasing pointing jitter. Recall that $\sigma_{\mathrm{e}}$ denotes the pointing-jitter standard deviation. At high SNR, the resulting ASE degradation (in nats/s/Hz) attributable solely to jitter can be obtained directly from (92) and is expressed as
@@
Remark 4.
The Figures 59 are grouped according to the previously discussed cases $A_{\mathrm{\scriptscriptstyle 0}}=1$ and $A_{\mathrm{\scriptscriptstyle 0}}<1$. Figures 5 and 6 correspond to turbulence-only conditions with $A_{\mathrm{\scriptscriptstyle 0}}=1$, whereas Figs. 7 and 8 consider composite turbulence and pointing-error conditions with the corresponding $A_{\mathrm{\scriptscriptstyle 0}}<1$ values listed in Table I. In Fig. 9, both channel cases are analyzed: the turbulence-only condition is characterized by $A_{\mathrm{\scriptscriptstyle 0}}=1$, while the turbulence with pointing errors is characterized by $A_{\mathrm{\scriptscriptstyle 0}}<1$.
Figure 5: Average spectral efficiency achievable at low SNRs for a GG turbulence channel without pointing jitter under adaptive coherent FSO transmission and synchronous heterodyne detection at the receiver (see Table I for the turbulence parameters).
@@
Figure 6: Probability density function (PDF) of the GG channel gain for the weak, moderate, and strong turbulence profiles described in Table I. Notice the increasing probability mass for higher channel gains with increasing turbulence strength.
@@
$\text{ASE pre-factor ratio}\,\approx\,\frac{a_{j}b_{j}}{a_{i}b_{i}}\,\cdot$ (105)